The length of a rectangle is 7 inches longer than it is wide. If the area is 170 square inches, what are the dimensions of the rectangle?
Answer by Cromlix(4381)

x = width x + 7 = length
x(x+7) = 170
x^2 + 7x - 170 = 0
(x + 17)( x - 10)
Therefore x = 10 as x = -17 no answer.
Width = 10 inches
Length = 17 inches